Subtract the complex numbers and simplify: 3+2i/1+2i-1/1-2i= Help! Thank you.
Found 2 solutions by venugopalramana, Prithwis:
Answer by venugopalramana(3286)

(3+2i)/(1+2i)-1/(1-2i)={(3+2i)(1-2i)/(1-2i)(1+2i)} -{1(1+2i)/(1+2i)(1-2i)}
={(3-6i+2i-4i^2)/(1-4i^2)} - {(1+2i)/(1-4i^2)}
={(7-4i)/5}- {(1+2i)/5}=(7-4i-1-2i)/5=(6-6i)/5=6(1-i)/5

Answer by Prithwis(166)

(3+2i)/(1+2i)] -1/(1-2i)
= (3+2i)(1-2i)-(1+2i)/[1^2 - (2i)^2]
= (-4i^2-4i+3-1-2i)/(1+4)
= 6(1-i)/5